GAgapito, Juan Carlo

Test a 400Hz

Rec_20090609_142233
F1:50%-50%.
F2:900-10.
Bin:1
CCD 39 Fre:400 Hz.
CCD 39 Read out:890
TT:11.00.
CCD47:3.5Hz
CCD47 Read out 2500

lambda/D=5 -> gain=0.4
lambda/D=7 -> gain=0.4
lambda/D=9 -> gain=0.5
lambda/D=11 -> gain=0.6

F1:50%-50%.
F2:900-10.
Bin:1
CCD 39 Fre:400 Hz.
CCD 39 Read out:890
CCD47:3.5Hz
CCD47 Read out 2500
-Lambda/D=11
Kalman: KF_21Hz_damp00_sigma1e6_compact_AR_142233_1000ph_400Hz
gain=0.01
gain=0.1
gain=0.3
gain=0.2 -> sembra meglio -> Optical Loop Diagnostic 4000 frames 20090724_105300.

NOTA: after 7 minutes in close loop the secondary starts to skip frames.

Try another close loop in the same conditions than before. The secondary starts skipping frame quickly. Saved loop optical loop diagnostics 20090724_110014 at OPEN LOOP.
It look like from the log that the actuator 228 is going out of range.
The WFS setup is with a REC for 351 Modes, at 400 Hz and modulation 11.
**Second try after Rest the secondary.-----the secondary starts to skip frame but the secondary shape is good and the loop can be keep close. Open the loop for safety.
saved optical loop diagnostic at close loop: 20090724_112247 during the skip frame is in red.
prova KF con 150 modi

gain = 0.01. ok

F1:50%-50%.
F2:900-10.
Bin:1
CCD 39 Fre:400 Hz.
CCD 39 Read out:890
TT:11.00.
CCD47:3.5Hz
CCD47 Read out 2500
gain = 0.2 -> 50 frames PSF 20090724/psf.fits -> Optical Loop Diagnostic 4000 frames 20090724_113545.
50 frames PSF 20090724/psf2.fits -> Optical Loop Diagnostic 4000 frames 20090724_114301.

gain= 0.5 -> 50 frames PSF 20090724/psf3.fits -> Optical Loop Diagnostic 4000 frames 20090724_114633.
50 frames PSF 20090724/psf4.fits -> Optical Loop Diagnostic 4000 frames 20090724_115336.

Test a 800 Hz


Rec_20090609_142233
F1:50%-50%.
F2:900-10.
Bin:1
CCD 39 Fre:800 Hz.
CCD 39 Read out:2500
TT:11.00.
CCD47:3.5Hz
CCD47 Read out 2500

OL -> 50 frames PSF 20090724/psf0.fits

lambda/D = 11 -> TT = 8.77
gain = 0.5
gain = 0.6
gain = 0.7

Kalman: KF_21Hz_damp00_sigma5e8_compact_AR_142233_1000ph_800Hz

gain = 0.01 (solo primi 150 modi) -> FunctionWarning - FunctionAllarm (orario circa 12:00)
si vede chiaramente un tilt! -> 10 frames PSF sequenza 20090724/psf_seq.fits -> Optical Loop Diagnostic 4000 frames 20090724_120754

lambda/D=12.5 -> TT = 9.6

gain = 0.01 (solo primi 150 modi)
si vede chiaramente un tilt! -> 10 frames PSF sequenza 20090724/psf_seq2.fits -> Optical Loop Diagnostic 4000 frames 20090724_121212

lambda/D=7.5 -> TT = 5.9

gain = 0.01 (solo primi 150 modi)
si vede chiaramente un tilt! -> 10 frames PSF sequenza 20090724/psf_seq3.fits -> Optical Loop Diagnostic 4000 frames 20090724_121354

lambda/D=5.0 -> TT=4.28

gain = 0.01 (solo primi 150 modi) si vede chiaramente un tilt! -> 10 frames PSF sequenza 20090724/psf_seq4.fits -> Optical Loop Diagnostic 4000 frames 20090724_121616

lambda/D=2.5

gain = 0.01 (solo primi 150 modi)
si vede chiaramente un tilt! -> 10 frames PSF sequenza 20090724/psf_seq5.fits -> Optical Loop Diagnostic 4000 frames 20090724_121828

CHARACTERISATION OF THE CCD47 WITH DIFFERENT SET UP FILTER & DICHROIC.

850-10900-10
50%-50%

Open Loop / close loop
Open Loop / close loop
90%-10% Open Loop / close loop Open Loop / close loop
600-1000 Open Loop / close loop Open Loop / close loop
700-1000 Open Loop / close loop Open Loop / close loop

SYSTEM SET UP:

Shape:
Rec:REc_20090609_142233
NModes:351
Gain: 0.5
CCD39:196 Hz.
CCD39 Read out.
CCD47: 3.5Hz.
CCD47 Read out:
FW1. changing in the measurements.
FW2. changing in the measurements.
TT:10.5.
pHOTON/SUBAPER:1000.

Number of frames saved ccd47 10 and sequence 10.

850-10900-10
50%-50%

psf1_ccd47 /psf2_ccd47
psf1_seq_ccd47/psf2_seq_ccd47
CL: 20090724_123531

psf3_ccd47 /psf4_ccd47
psf3_seq_ccd47/psf4_seq_ccd47
CL: 20090724_124219
90%-10% psf5_ccd47/ psf6_ccd47
psf5_seq_ccd47/psf?_seq_ccd47
CL: 20090724_125842.
psf7_ccd47 / psf8_ccd47
psf7_seq_ccd47 / psf7_seq_ccd47
CL: 20090724_130347.
600-1000 psf10_ccd47 / psf9_ccd47
--- / psf9_seq_ccd47
CL: 20090724_154214
psf12_ccd47 / psf11_ccd47
psf12_seq_ccd47 / psf11_seq_ccd47
CL: 20090724_162300
700-1000 psf13_ccd47 / psf14_ccd47
psf_seq_ccd47 / psf14_seq_ccd47
CL: 20090724_164325
psf15_ccd47 / psf16_ccd47
------ / psf16_seq_ccd47
CL: 20090724_165505

NOTA: the secondary went to skipping frame: time12:49:00, the recover skip frame doen't work, it needed to RIP the secondary using the recovery failE BUTTON.

OPEN LOOP DIAGNOSTICS SAVED TO DETERMINE THE OSCILLATION FREQUENCY of the TOWER.

External wind ??? calm
20090724_152409, 20090724_152648

Gain:0
CCD39:200Hz.
TT:10.

20090724_153129, 20090724_153319
Gain:0
CCD39:200Hz.
TT:5

20090724_170722 , 20090724_170837
gain:0
CCD39: 400Hz.
TT:2.


In pausa pranzo Doug ha lavorato sulla IRTC


Kalman: KF_21Hz_damp00_sigma1e8_compact_AR_142233_1000ph_800Hz

lambda/D=12.5 -> TT = 9.6

gain = 0.01 (solo primi 150 modi)
si vede chiaramente un tilt! -> 10 frames PSF sequenza 20090724/psf_seq6.fits -> Optical Loop Diagnostic 4000 frames 20090724_141247

lambda/D=7.5 -> TT = 5.9

gain = 0.01 (solo primi 150 modi)
si vede chiaramente un tilt! -> 10 frames PSF sequenza 20090724/psf_seq7.fits -> Optical Loop Diagnostic 4000 frames 20090724_141524

lambda/D=5.0 -> TT=4.28

gain = 0.01 (solo primi 150 modi)
si vede chiaramente un tilt! -> 10 frames PSF sequenza 20090724/psf_seq8.fits -> Optical Loop Diagnostic 4000 frames 20090724_141715

lambda/D=2.5

gain = 0.01 (solo primi 150 modi)
si vede chiaramente un tilt! -> 10 frames PSF sequenza 20090724/psf_seq9.fits -> Optical Loop Diagnostic 4000 frames 20090724_141906

lambda/D=1.5

gain = 0.01 (solo primi 150 modi)
si vede chiaramente un tilt! -> 10 frames PSF sequenza 20090724/psf_seq10.fits -> Optical Loop Diagnostic 4000 frames 20090724_142216

TEST KF vs INT

KF

FW1 - DC 600/1000
FW2 - 900/10

fc = 400Hz
lambda/D = 5
Kalman: KF_21Hz_damp00_sigma1e6_compact_AR_142233_1000ph_400Hz

solo primi 150 modi -> chiude a 0.01, 0.2, 0.3, 0.5 (si vede un modo alto residuo)

- prova con integratore:

dist_atmo_08_1000Hz_15ms_s35_80pix.fits

non si vede niente sul ccd47 -> mettiamo 50/50 sulla FW1 -> anche con 0.5 non riesce a correggere bene

dist_atmo_08_1000Hz_15ms_s35.fits

anche con 0.5 non riesce a correggere bene

- prova con KF:

con gain = 0.1 funziona

con gain = 0.3 skip del secondario (recover skip non ha funzionato 2 volte) -> attuatori 131, 149, 481

dist_atmo_08_1000Hz_15ms_s35_80pix.fits

con gain = 0.3 - 0 per KF

INT

FW1 600/1000
FW2 850/10
fc = 400Hz
lambda/D =2.5
fotoni = 950

Rec_20090609_142233

CL gain = 0.5 senza disturbo -> 4000/50 frame -> 154724
OL gain = 0 senza disturbo -> 4000/50 frame -> 154856

dist_atmo_08_1000Hz_15ms_s35_80pix.fits

OL gain = 0 con disturbo -> 4000/50 frame -> 155008
CL gain = 0.5 con disturbo -> 4000/50 frame -> 155117

ERRORE -> disturb_f21.0000Hz_amp10.0000mas_tt0_b0_4p_dist_atmo_08_1000Hz_15ms_s35_80pix_fc400.000Hz.fits

OL gain = 0 con disturbo -> 4000/50 frame -> 155531
CL gain = 0.5 con disturbo -> 4000/50 frame -> 155635

ERRORE -> disturb_f21.0000Hz_amp20.0000mas_tt0_b0_4p_dist_atmo_08_1000Hz_15ms_s35_80pix_fc400.000Hz.fits

OL gain = 0 con disturbo -> 4000/50 frame -> 155824
155933 errore (CL senza disturbo)
CL gain = 0.5 con disturbo -> 4000/50 frame -> 160029

ERRORE -> disturb_f21.0000Hz_amp50.0000mas_tt0_b0_4p_dist_atmo_08_1000Hz_15ms_s35_80pix_fc400.000Hz.fits

OL gain = 0 con disturbo -> 4000/50 frame -> 160314
CL gain = 0.5 con disturbo -> 4000/50 frame -> 160416

ERRORE -> disturb_f21.0000Hz_amp100.0000mas_tt0_b0_4p_dist_atmo_08_1000Hz_15ms_s35_80pix_fc400.000Hz.fits

OL gain = 0 con disturbo -> 4000/50 frame -> 160725
CL gain = 0.5 con disturbo -> 4000/50 frame -> 160903

ERRORE -> disturb_f21.0000Hz_amp150.0000mas_tt0_b0_4p_dist_atmo_08_1000Hz_15ms_s35_80pix_fc400.000Hz.fits

OL gain = 0 con disturbo -> 4000/50 frame -> 161134
CL gain = 0.5 con disturbo -> 4000/50 frame -> 161316

disturb_f21.0000Hz_amp150.0000mas_tt0_b0_4p_dist_atmo_08_1000Hz_15ms_s35_80pix_fc400.000Hz.fits

NOTA:ALL THE DISTURBANCE FILES GENERATED FROM THE KALMAN ARE NOT APPLIED TO THE SECONDARY AND ALL THE MEASUREMENTS HAVE BEEN TAKEN WITH THE DEFAULT DISTURBANCES. LOOKING THE LOG FILES LOOKS LIKE THAT THE SYSTEM DIDN'T ACCEPT THE DISTURBANCE FILE DUE TO A WRONG PARAMETERS
THE BUG HAS BEEN NOTIFIED TO ALFIO, AND HE HAS FIXED.

-- GuidoAgapito - 2009-07-24
Topic revision: r4 - 27 Jul 2009, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback